OpenStudy (anonymous):

prove that the function f(z) = Re(z) does not have a derivative for any value of z

OpenStudy (anonymous):

@phi @Hero @jhonyy9

OpenStudy (anonymous):

Use the limit definition of the derivative and you have it approaching different values along different lines

OpenStudy (anonymous):

how to prove it using philosophy of course very deeply??

OpenStudy (sirm3d):

use Cauchy-Riemann conditions for differentiability. \[\newcommand {pd}{\partial}\frac{\pd u}{\pd x}=\frac{\pd v}{\pd y},\qquad \frac{\pd u}{\pd y}=-\frac{\pd v}{\pd x}\] given \[f(z)=u(x,y) + v(x,y)i\]

OpenStudy (sirm3d):

let \[z=x+yi\] \[f(z)= \text{Re } z=x+0i\]

OpenStudy (sirm3d):

\[u_x=1,v_y=0\] the first condition is not satisfied.

OpenStudy (anonymous):

@sirm3d did you know symmetric tops of molecule??